OpenStudy (anonymous):

Definite integral question Use the area of the integrand to evaluate the integral. Question is inside the post.

OpenStudy (anonymous):

\[\int\limits_{\pi}^{2\pi}\theta d \theta\]

OpenStudy (anonymous):

The answer is \[3\pi^2 /2 \] but I just have not idea how.

OpenStudy (anonymous):

Do you know what an anti-derivative is?